Chris has bought 350 feet of fencing to build a rectangular pen. Using your answer in the previous question, which of the following represents area A in terms of the width w?
THE ANSWER IN PREVIOUS QUESTION IS L=175-W

Answer by Boreal(15235)

the half-perimeter is L+W=175 feet. That is why L=175-W
area is L*W=W(175-W)=175W-W^2
Draw this out and label, and it becomes a lot more evident.